# Transport properties in a model of confined granular mixtures at moderate densities

### v1

@@ VI Granular binary mixtures
Here, $D_{11}^{*}(1)$, $D_{12}^{*}(1)$, and $D_{1}^{*T}(1)$ refer to the values of the diffusion coefficients for elastic collisions ($\alpha=1$). Although the deviations of the diffusion transport coefficients from their forms for elastic collisions in the $\Delta$-model are generally smaller than those in the conventional IHS model,Garzó and Dufty (2002); Garzó, Montanero, and Dufty (2006); Garzó and Montanero (2007) we observe that the effect of inelasticity on mass transport in two-dimensional, confined granular mixtures can be significant, especially for mixtures of particles with the same mass density (the case $\sigma_{1}/\sigma_{2}=3$ and $m_{1}/m_{2}=9$). Regarding the dependence on the mass ratio, we find that the (dimensionless) coefficients $D_{11}^{*}$ and $D_{12}^{*}$ increase monotonically with respect to their elastic forms with decreasing $\alpha$ (i.e, as the inelasticity in collisions increases), except for $D_{12}^{*}$ when $m_{1}/m_{2}=(\sigma_{1}/\sigma_{2})^{2}=9$, where this coefficient exhibits weak non-monotonic dependence on $\alpha$ for relatively high inelasticity.
@@ VI.3 Shear and bulk viscosities
From Eqs. (67)–(69), it is clear that evaluating the bulk viscosity $\eta_{b}$ requires knowing the first-order contributions $\varpi_{i}$ to the partial temperatures. However, determining the above coefficients is beyond the scope of this paper, as it involves lengthy and complex calculations. On the other hand, as mentioned in section V, according to the previous results derived from the IHS modelGómez González and Garzó (2019) and for low-density confined granular mixturesGarzó, Brito, and Soto (2018) the influence of $\varpi_{i}$ on the value of the bulk viscosity $\eta_{b}$ is generally very small. Thus, for practical purposes, the bulk viscosity $\eta_{b}$ can be well estimated by $\eta_{b}^{(\text{I})}$. The ratio $\eta_{b}^{*}(\alpha)/\eta_{b}^{*}(1)$ is plotted in Fig. 8 as a function of the (common) coefficient of restitution $\alpha$ for $d=2$, $x_{1}=0.5$, $\phi=0.2$ and three different mixtures. Here, $\eta_{b}^{*}=(\nu/nT)\eta_{b}(1)$ where $\eta_{b}^{*}(1)$ is the (dimensionless) bulk viscosity for elastic collisions. As with the case of the shear viscosity, we observe that collisional dissipation has a weaker influence on the bulk viscosity than in the IHS model.Gómez González and Garzó (2019)
@@ VII An application: Thermal diffusion segregation in a confined granular dense mixture
$\Lambda\frac{\partial\ln T}{\partial y}=\frac{\partial}{\partial y}\ln\left(\frac{n_{1}}{n_{2}}\right).$ (119)
A binary mixture has been considered here, and, for simplicity, it is assumed that, in two-dimensional systems, gradients occur only along the y-axis. Furthermore, without loss of generality, we assume that $\sigma_{1}>\sigma_{2}$ and that the gravitational field is parallel to the thermal gradient, so $\mathbf{g}=-g\hat{e}_{y}$ where $\hat{e}_{y}$ is the unit vector in the positive direction of the y-axis This means that the bottom plate is hotter than the top plate, so $\partial_{y}T<0$ (see for instance the sketch of Fig. 3 of Ref. Garzó, Brito, and Soto, 2024a).
According to Eq. (119), when $\Lambda>0$, the larger particles 1 tend to accumulate near the top (cold) plate, since $\partial_{y}\ln(n_{1}/n_{2})>0$. Conversely, when $\Lambda<0$, the larger particles 1 tend to accumulate near the bottom (hot) plate, since $\partial_{y}\ln(n_{1}/n_{2})<0$. Now, we write the thermal diffusion factor $\Lambda$ in terms of the (dimensionless) diffusion transport coefficients $D_{11}^{*}$, $D_{12}^{*}$, and $D_{1}^{*T}$. First, the momentum balance equation (12) leads to
$\frac{\partial p}{\partial y}=-\rho g,$ (120)
where we have accounted for that $P_{ij}=p\delta_{ij}$. According to Eq. (51), the hydrostatic pressure $p=nTp^{*}$ where the dimensionless pressure $p^{*}$ depends on $y$ through its dependence on $x_{1}$, $\phi$, and $\Delta^{*}=\Delta/v_{\text{th}}(T(y))$. Thus, in dimensionless form, Eq. (120) can be rewritten as
@@
$\Lambda_{1}=\frac{\partial_{y}\ln n_{1}}{\partial_{y}\ln T},\quad\Lambda_{2}=\frac{\partial_{y}\ln n_{2}}{\partial_{y}\ln T},$ (122)
@@
$g^{*}=\frac{\rho g}{n\left(\frac{\partial T}{\partial y}\right)}<0$ (125)
@@
The explicit form of $\Lambda$ on the parameters of the mixture can be obtained when one substitutes Eqs. (170), (172) and (173) for the diffusion transport coefficients $D_{1}^{T*}$, $D_{11}^{*}$, and $D_{12}^{*}$, respectively, and Eq. (51) for $p^{*}$ (and its corresponding derivatives $\xi_{i}$) into Eq. (129). This yields the dependence of the thermal diffusion factor on the parameter space of the problem, including mass and size ratios, mole fraction, scaled gravity, solid volume fraction, and coefficients of restitution. In particular, inspecting the dependence of the denominator $x_{1}x_{2}(\xi_{2}D_{11}^{*}-\xi_{1}D_{12}^{*})$ of Eq. (129) on the parameters of the mixture shows that it is usually positive. Thus, the condition of setting the thermal diffusion factor to zero yields the curves delineating the regimes between positive and negative values of the thermal diffusion factor. This yields the segregation criterion
@@ VII.3 Tracer limit for moderate densities
Figure 10: Plot of the dependence of the marginal segregation curve ($\Lambda=0$) on the coefficient of restitution $\alpha_{11}$ in for a two-dimensional system with $x_{1}=0.5$, $\sigma_{1}/\sigma_{2}=2$, $\alpha_{22}=0.9$, $\alpha_{12}=(\alpha_{11}+\alpha_{12})/2$, and three different values of the solid volume fraction: $\phi=0$ (a); $\phi=0.1$ (b); and $\phi=0.2$ (c). The limiting case $|g^{*}|\to\infty$ is considered.
@@ VII.4 Moderately dense regime in a confined granular binary mixture
We now consider granular binary mixtures with an arbitrary concentration at moderate densities. As in Ref. Garzó, Brito, and Soto, 2024a, although gravity and the thermal gradient are of the same order of magnitude in our analysis, it is interesting for illustrative purposes to separate the influence of each term in Eq. (130) on segregation. Thus, we first consider cases where gravity is absent ($g=0$ but $\partial_{y}T\neq 0$) or thermalized systems where the effect of gravity on segregation is much more important than that of thermal gradient ($g\neq 0$ but $\partial_{y}T\to 0$).
@@
A phase diagram for the marginal segregation curve ($\Lambda=0$) in the $(\sigma_{1}/\sigma_{2};m_{1}/m_{2})$-plane for a two-dimensional system is plotted in Fig. 9. We have considered mixtures with a concentration $x_{1}=0.5$, a (common) coefficient of restitution $\alpha=0.5$, and three different values of the solid volume fraction: a dilute mixture ($\phi=0$) and two mixtures with moderate densities ($\phi=0.1$ and $\phi=0.2$). In general, we observe that density primarily increases weakly the size of the region corresponding to $\Lambda<0$. This means that larger particles tend to accumulate near the hot plate as the system becomes denser. Additionally, at a given diameter ratio, the region with larger particles attempting to move toward the cold plate (i.e., the region $\Lambda>0$) appears essentially when larger particles are much heavier than the other species. Comparing these results with those obtained in the conventional IHS model shows qualitative agreement between both models (see for instance, Fig. 5 of Ref. Garzó, 2011).
VII.4.2 Thermalized systems ($\partial_{y}T\to 0$)
@@
To illustrate this limiting segregation case ($|g^{*}|\to\infty$), Fig. 10 shows the dependence of the marginal segregation curve on the coefficient of restitution $\alpha_{11}$ for a two-dimensional system. The system has the following parameters: $x_{1}=0.5$, $\sigma_{1}/\sigma_{2}=2$, $\alpha_{22}=0.9$, $\alpha_{12}=(\alpha_{11}+\alpha_{22})/2$, and three values of the solid volume fraction ($\phi=0$, 0.1 and 0.2). For a given value of the coefficient of restitution $\alpha_{11}$, we observe that the thermal diffusion factor $\Lambda$ is always negative when the larger particles are much heavier than the smaller ones. In this case, the larger particles accumulate near the hot plate. The effect of density on the marginal segregation curve is also apparent, as the region where the thermal diffusion factor becomes positive (and larger particles move toward the cold plate) increases with increasing density.
@@ VII.4.3 General case
Finally, we consider the general case for finite values of reduced gravity, $|g^{*}|$. To illustrate this situation, Fig. (11) plots the marginal segregation curve ($\Lambda=0$) versus the common coefficient of restitution $\alpha_{ij}\equiv\alpha$ for a two-dimensional system with $x_{1}=0.5$, $\sigma_{1}/\sigma_{2}=2$, and three values of the solid volume fraction ($\phi=0$, 0.1 and 0.2). As can be seen, the behavior of the marginal segregation curve with inelasticity is quite similar to that found when $|g^{*}|\to\infty$. At a given value of $\alpha$, the region $\Lambda>0$ increases as the density of the system increases.
To the best of our knowledge, there are no computer simulations available in the granular literature for comparison with the theoretical results reported here for the marginal segregation curve ($\Lambda=0$) in the context of the $\Delta$-model for mixtures. We expect the current results to encourage the development of such simulations, which will help assess the reliability of the theoretical results derived for the thermal diffusion factor $\Lambda$.
@@ VIII Discussion
The objective of the present study is to extend the results derived in Refs. Garzó, Dufty, and Hrenya, 2007; Garzó, Hrenya, and Dufty, 2007 to confined quasi-two-dimensional granular mixtures. In this geometry, the system is confined in a box in which the $z$-direction is slightly larger than one particle diameter. The box is vertically vibrated, causing the particles to gain energy through collisions with the walls. This energy gained is then transferred to the horizontal degrees of freedom of grains when collisions between particles take place.
However, due to the technical difficulties associated with describing this type of system using kinetic theory, a coarse-grained model is typically employed. In this collisional model (referred to as the $\Delta$-model)Brito, Risso, and Soto (2013), particles move only in two dimensions. When particles collide, part of the gained energy (due to their collisions with the vibrating walls) is released into their horizontal degrees of freedom. The model accounts for this effect by introducing positive factors, denoted by $\Delta_{ij}$, which are added to the relative motion of colliding spheres. The quantities $\Delta_{ij}$ can be related to the intensity of vertical vibrations in experiments.Maynar, García de Soria, and Brey (2019)
@@
To achieve analytical expressions for the transport coefficients, the relevant state of a two-dimensional, confined mixture with a stationary temperature has been considered. The explicit forms for the diffusion coefficients $D_{ij}$ and $D_{i}^{T}$ and the kinetic shear viscosity $\eta_{k}$ have been obtained by taking the leading terms in a Sonine polynomial expansion of the first-order distribution functions. This is the standard procedure for determining these coefficients for elasticChapman and Cowling (1970); Ferziger and Kaper (1972) and inelasticGarzó (2019) mixtures. In the case of the bulk viscosity $\eta_{b}$, the contributions coming from the first-order partial temperatures $\varpi_{i}$ have been neglected. This approximation is based on the results obtained in the IHS modelGómez González and Garzó (2019) where the influence of $\varpi_{i}$ on $\eta_{b}$ has been in general shown quite small. We expect that this approximation is accurate in the $\Delta$-model as well. Under the above approximations, for binary mixtures, the diffusion coefficients are given by Eqs. (170), (172), and (173), the shear viscosity by Eqs. (81) and (118), and the bulk viscosity by Eq. (80).
@@
We have also analyzed thermal diffusion segregation induced by both gravity and a thermal gradient as an application of the derived kinetic theory. In this situation, the so-called thermal diffusion factor $\Lambda$ [defined by Eq. (119)] provides a segregation criterion. An explicit expression of $\Lambda$ in terms of the (dimensionless) diffusion coefficients $D_{11}^{*}$, $D_{12}^{*}$, and $D_{1}^{T*}$, the (reduced) pressure $p^{*}$ and the (reduced) gravity $g^{*}$ is given by Eq. (129). Assuming the lower plate of the container is hotter than the top plate, when the thermal diffusion factor is positive ($\Lambda>0$), the larger particles tend to accumulate near the cold plate against gravity, while the smaller particles sink to the bottom plate. When the thermal diffusion factor is negative ($\Lambda<0$), the opposite effect occurs. The condition $\Lambda=0$ provides the focus line between the two opposite behaviors. Comparing these results with those obtained in the low-density regime Garzó, Brito, and Soto (2024a, b) shows that the effect of density on the phase diagrams for the marginal segregation curve ($\Lambda=0$) is generally significant. Specifically, the region $\Lambda<0$ is increased as the system becomes more dense in the absence of gravity (see Fig. 9), but decreases in the presence of gravity (see Figs. 10 and 11).
@@
As said before, one of our future plans within the $\Delta$-model is to determine the heat flux transport coefficients for dense mixtures. Knowing the complete set of Navier-Stokes transport coefficients will enable us to perform a linear stability analysis of the homogeneous steady state, among other applications. In particular, since the homogeneous steady state is stable in the dilute limit Garzó, Brito, and Soto (2021), we want to see if density corrections to the transport coefficients can modify the stability of the homogeneous state. Additionally, the reliability of the theoretical results derived here, which were obtained under certain approximations, should be assessed against computer simulations. One possible project is to carry out computer simulations to measure shear viscosity. As for dry granular mixtures, Montanero and Garzó (2003); Garzó and Montanero (2003); Chamorro and Garzó (2023) we plan to perform simulations on a granular mixture subjected to simple shear flow, where the cooling effects associated with collision dissipation are precisely offset by the presence of the $\Delta$ parameters in the collisional rules. Work along these lines is in progress.

### v2

@@ VI Granular binary mixtures
In contrast to the monocomponent limiting case, for which the dependence of $\Delta^{*}$ on $\alpha$ is explicitly given by Eq. (111), the dependence of $\Delta^{*}$ and $T_{1}^{(0)}/T_{2}^{(0)}$ on the parameters of the mixture is determined by numerically solving the set of coupled equations $\zeta_{1}^{*}=0$ and $\zeta_{2}^{*}=0$. An estimate of the partial cooling rates $\zeta_{i}^{*}$ by assuming Maxwellian distributions for the zeroth-order solutions $f_{i}^{(0)}(\mathbf{V})$ are given by Eq. (52). As previously mentioned, the forms (52) yield good agreement with computer simulations for both $\Delta^{*}$ and $T_{1}^{(0)}/T_{2}^{(0)}$.Brito, Soto, and Garzó (2020)
@@
Here, $D_{11}^{*}(1)$, $D_{12}^{*}(1)$, and $D_{1}^{*T}(1)$ refer to the values of the diffusion coefficients for elastic collisions ($\alpha=1$). It must be recalled that $\Delta^{*}$ and $T_{1}^{(0)}/T_{2}^{(0)}$ change in Fig. 3 with the parameters of the mixture. Although the deviations of the diffusion transport coefficients from their forms for elastic collisions in the $\Delta$-model are generally smaller than those in the conventional IHS model,Garzó and Dufty (2002); Garzó, Montanero, and Dufty (2006); Garzó and Montanero (2007) we observe that the effect of inelasticity on mass transport in two-dimensional, confined granular mixtures can be significant, especially for mixtures of particles with the same mass density (the case $\sigma_{1}/\sigma_{2}=3$ and $m_{1}/m_{2}=9$). Regarding the dependence on the mass ratio, we find that the (dimensionless) coefficients $D_{11}^{*}$ and $D_{12}^{*}$ increase monotonically with respect to their elastic forms with decreasing $\alpha$ (i.e, as the inelasticity in collisions increases), except for $D_{12}^{*}$ when $m_{1}/m_{2}=(\sigma_{1}/\sigma_{2})^{2}=9$, where this coefficient exhibits weak non-monotonic dependence on $\alpha$ for relatively high inelasticity.
@@ VI.3 Shear and bulk viscosities
From Eqs. (69)–(71), it is clear that evaluating the bulk viscosity $\eta_{b}$ requires knowing the first-order contributions $\varpi_{i}$ to the partial temperatures. However, determining the above coefficients is beyond the scope of this paper, as it involves lengthy and complex calculations. On the other hand, as mentioned in section V, according to the previous results derived from the IHS modelGómez González and Garzó (2019) and for low-density confined granular mixturesGarzó, Brito, and Soto (2018) the influence of $\varpi_{i}$ on the value of the bulk viscosity $\eta_{b}$ is generally very small. Thus, for practical purposes, the bulk viscosity $\eta_{b}$ can be well estimated by $\eta_{b}^{(\text{I})}$. The ratio $\eta_{b}^{*}(\alpha)/\eta_{b}^{*}(1)$ is plotted in Fig. 9 as a function of the (common) coefficient of restitution $\alpha$ for $d=2$, $x_{1}=0.5$, $\phi=0.2$ and three different mixtures. Here, $\eta_{b}^{*}=(\nu/nT)\eta_{b}$ where $\eta_{b}^{*}(1)$ is the (dimensionless) bulk viscosity for elastic collisions. As with the case of the shear viscosity, we observe that collisional dissipation has a weaker influence on the bulk viscosity than in the IHS model.Gómez González and Garzó (2019)
@@ VII An application: Thermal diffusion segregation in a confined granular dense mixture
Figure 10: Illustration of the segregation process behavior in a granular binary mixture. The BNE effect ($\Lambda>0$) occurs when the large (brown) particles tend to accumulate near the cold (top) plate of the system. The RBNE effect ($\Lambda<0$) occurs when the large (brown) particles tend to accumulate near the hot (bottom) plate of the system.
@@
$\Lambda\frac{\partial\ln T}{\partial z}=\frac{\partial}{\partial z}\ln\left(\frac{n_{1}}{n_{2}}\right).$ (122)
A binary mixture has been considered here, and, for simplicity, it is assumed that, in two-dimensional systems, gradients occur only along the $z$-axis. Furthermore, without loss of generality, we assume that $\sigma_{1}>\sigma_{2}$ and that the gravitational field is parallel to the thermal gradient, so $\mathbf{g}=-g\hat{e}_{z}$ where $\hat{e}_{z}$ is the unit vector in the positive direction of the $z$-axis This means that the bottom plate is hotter than the top plate, so $\partial_{z}T<0$ (see for instance the sketch of Fig. 10).
According to Eq. (122), when $\Lambda>0$, the larger particles 1 rise to the top (cold) plate, since $\partial_{z}\ln(n_{1}/n_{2})>0$. We recall that this effect is usually known as the Brazil nut effect (BNE). Conversely, when $\Lambda<0$, the larger particles 1 sink to the bottom (hot) plate, since $\partial_{z}\ln(n_{1}/n_{2})<0$. This effect is known as the reverse Brazil nut effect (RBNE). Now, we write the thermal diffusion factor $\Lambda$ in terms of the (dimensionless) diffusion transport coefficients $D_{11}^{*}$, $D_{12}^{*}$, and $D_{1}^{*T}$. First, the momentum balance equation (12) leads to
$\frac{\partial p}{\partial z}=-\rho g,$ (123)
where we have accounted for that $P_{ij}=p\delta_{ij}$. According to Eq. (53), the hydrostatic pressure $p=nTp^{*}$ where the dimensionless pressure $p^{*}$ depends on $z$ through its dependence on $x_{1}$, $\phi$, and $\Delta^{*}=\Delta/v_{\text{th}}(T(z))$. Thus, in dimensionless form, Eq. (123) can be rewritten as
@@
$\Lambda_{1}=\frac{\partial_{z}\ln n_{1}}{\partial_{z}\ln T},\quad\Lambda_{2}=\frac{\partial_{z}\ln n_{2}}{\partial_{z}\ln T},$ (125)
@@
$g^{*}=\frac{\rho g}{n\left(\frac{\partial T}{\partial z}\right)}<0$ (128)
@@
The explicit form of $\Lambda$ on the parameters of the mixture can be obtained when one substitutes Eqs. (173), (175) and (176) for the diffusion transport coefficients $D_{1}^{T*}$, $D_{11}^{*}$, and $D_{12}^{*}$, respectively, and Eq. (53) for $p^{*}$ (and its corresponding derivatives $\xi_{i}$) into Eq. (132). This yields the dependence of the thermal diffusion factor on the parameter space of the problem, including mass and size ratios, mole fraction, scaled gravity, solid volume fraction, and coefficients of restitution. In particular, inspecting the dependence of the denominator $x_{1}x_{2}(\xi_{2}D_{11}^{*}-\xi_{1}D_{12}^{*})$ of Eq. (132) on the parameters of the mixture shows that it is usually positive. Thus, the condition of setting the thermal diffusion factor to zero yields the curves delineating the regimes between the segregation toward the cold and the hot wall (BNE/RBNE transition). This yields the segregation criterion
@@ VII.3 Tracer limit for moderate densities
Figure 12: Plot of the dependence of the marginal segregation curve ($\Lambda=0$) on the coefficient of restitution $\alpha_{11}$ in for a two-dimensional system with $x_{1}=0.5$, $\sigma_{1}/\sigma_{2}=2$, $\alpha_{22}=0.9$, $\alpha_{12}=(\alpha_{11}+\alpha_{22})/2$, and three different values of the solid volume fraction: $\phi=0$ (a); $\phi=0.1$ (b); and $\phi=0.2$ (c). The limiting case $|g^{*}|\to\infty$ is considered.
@@ VII.4 Moderately dense regime in a confined granular binary mixture
We now consider granular binary mixtures with an arbitrary concentration at moderate densities. As in Ref. Garzó, Brito, and Soto, 2024a, although gravity and the thermal gradient are of the same order of magnitude in our analysis, it is interesting for illustrative purposes to separate the influence of each term in Eq. (133) on segregation. Thus, we first consider cases where gravity is absent ($g=0$ but $\partial_{z}T\neq 0$) or thermalized systems where the effect of gravity on segregation is much more important than that of thermal gradient ($g\neq 0$ but $\partial_{z}T\to 0$).
@@
A phase diagram delineating the regimes between BNE and RBNE in the $(\sigma_{1}/\sigma_{2};m_{1}/m_{2})$-plane for a two-dimensional system is plotted in Fig. 11. The lines are determined from the condition $\Lambda=0$. We have considered mixtures with a concentration $x_{1}=0.5$, a (common) coefficient of restitution $\alpha=0.5$, and three different values of the solid volume fraction: a dilute mixture ($\phi=0$) and two mixtures with moderate densities ($\phi=0.1$ and $\phi=0.2$). In general, we observe that density primarily increases weakly the size of the RBNE region ($\Lambda<0$). This means that larger particles tend to accumulate near the hot plate as the system becomes denser. Additionally, at a given diameter ratio, the region with larger particles attempting to move toward the cold plate (i.e., the BNE region where $\Lambda>0$) appears essentially when larger particles are much heavier than the other species. Comparing these results with those obtained in the conventional IHS model shows qualitative agreement between both models (see for instance, Fig. 5 of Ref. Garzó, 2011).
VII.4.2 Thermalized systems ($\partial_{z}T\to 0$)
@@
To illustrate this limiting segregation case ($|g^{*}|\to\infty$), Fig. 12 shows the dependence of the marginal segregation curve on the coefficient of restitution $\alpha_{11}$ for a two-dimensional system. The system has the following parameters: $x_{1}=0.5$, $\sigma_{1}/\sigma_{2}=2$, $\alpha_{22}=0.9$, $\alpha_{12}=(\alpha_{11}+\alpha_{22})/2$, and three values of the solid volume fraction ($\phi=0$, 0.1 and 0.2). For a given value of the coefficient of restitution $\alpha_{11}$, we observe that the thermal diffusion factor $\Lambda$ is always negative (RBNE effect) when the larger particles are much heavier than the smaller ones. In this case, the larger particles accumulate near the hot plate. The effect of density on the marginal segregation curve is also apparent, as the region where the thermal diffusion factor becomes positive (and larger particles move toward the cold plate, BNE effect) increases with increasing density.
@@ VII.4.3 General case
Finally, we consider the general case for finite values of reduced gravity, $|g^{*}|$. To illustrate this situation, Fig. (13) plots the marginal segregation curve ($\Lambda=0$) versus the common coefficient of restitution $\alpha_{ij}\equiv\alpha$ for a two-dimensional system with $x_{1}=0.5$, $\sigma_{1}/\sigma_{2}=2$, and three values of the solid volume fraction ($\phi=0$, 0.1 and 0.2). As can be seen, the behavior of the marginal segregation curve with inelasticity is quite similar to that found when $|g^{*}|\to\infty$. At a given value of $\alpha$, the size of BNE region ($\Lambda>0$) increases as the density of the system increases.
To the best of our knowledge, there are no computer simulations available in the granular literature for comparison with the theoretical results reported here for the BNE/RBNE transition in the context of the $\Delta$-model for mixtures. We expect the current results to encourage the development of such simulations, which will help assess the reliability of the theoretical results derived for the thermal diffusion factor $\Lambda$.
@@ VIII Discussion
The objective of the present study is to extend the results derived in Refs. Garzó, Dufty, and Hrenya, 2007; Garzó, Hrenya, and Dufty, 2007 for moderately dense granular mixtures in the context of the conventional IHS model to the $\Delta$-model.Brito, Risso, and Soto (2013) The $\Delta$-model is a coarse-grained collisional model that attempts to incorporate collisional energy injection into the dynamics of granular particles in a minimal way, particularly in situations that mimic vertically vibrated and confined systems. As illustrated in Fig. 1, in this geometry the system is confined in a box in which the $z$-direction is slight larger than one particle diameter. The box is vertically vibrated, causing the particles to gain energy through collisions with the walls. This energy gained is then transferred to the horizontal degrees of freedom of grains when collisions between particles take place.
Due to the technical difficulties associated with describing this type of system using kinetic theory, the $\Delta$-model is usually considered as the starting point to determine the dynamic properties of the system. In the $\Delta$-model, when particles collide, part of the gained energy (due to their collisions with the vibrating walls) is released into their horizontal degrees of freedom. The model accounts for this effect by introducing positive factors, denoted by $\Delta_{ij}$, which are added to the relative motion of colliding spheres. The quantities $\Delta_{ij}$ can be related to the intensity of vertical vibrations in experiments.Maynar, García de Soria, and Brey (2019)
@@
To achieve analytical expressions for the transport coefficients, the relevant state of a confined mixture with a stationary temperature has been considered. Additionally, although the analytical results have been derived in this paper for a $d$-dimensional system ($d=2$ for hard disks and $d=3$ for hard sphere), the case $d=2$ has been specifically considered. This is because the $\Delta$-model was primarily proposed to reproduce the results obtained in confined, quasi-two-dimensional setups. We expect the main conclusions reported here for $d=2$ to be similar to those for $d=3$.
The explicit forms for the diffusion coefficients $D_{ij}$ and $D_{i}^{T}$ and the kinetic shear viscosity $\eta_{k}$ have been obtained by taking the leading terms in a Sonine polynomial expansion of the first-order distribution functions. This is the standard procedure for determining these coefficients for elasticChapman and Cowling (1970); Ferziger and Kaper (1972) and inelasticGarzó (2019) mixtures. In the case of the bulk viscosity $\eta_{b}$, the contributions coming from the first-order partial temperatures $\varpi_{i}$ have been neglected. This approximation is based on the results obtained in the IHS modelGómez González and Garzó (2019) where the influence of $\varpi_{i}$ on $\eta_{b}$ has been in general shown quite small. We expect that this approximation is accurate in the $\Delta$-model as well. Under the above approximations, for binary mixtures, the diffusion coefficients are given by Eqs. (173), (175), and (176), the shear viscosity by Eqs. (82) and (121), and the bulk viscosity by Eq. (81).
@@
We have also analyzed thermal diffusion segregation induced by both gravity and a thermal gradient as an application of the derived kinetic theory. In this situation, the so-called thermal diffusion factor $\Lambda$ [defined by Eq. (122)] provides a segregation criterion. An explicit expression of $\Lambda$ in terms of the (dimensionless) diffusion coefficients $D_{11}^{*}$, $D_{12}^{*}$, and $D_{1}^{T*}$, the (reduced) pressure $p^{*}$ and the (reduced) gravity $g^{*}$ is given by Eq. (132). Assuming the lower plate of the container is hotter than the top plate, when the thermal diffusion factor is positive ($\Lambda>0$, BNE effect), the larger particles tend to accumulate near the cold plate against gravity, while the smaller particles sink to the bottom plate. When the thermal diffusion factor is negative ($\Lambda<0$, RBNE effect), the opposite effect occurs. The condition $\Lambda=0$ provides the focus line between the two opposite behaviors. Comparing these results with those obtained in the low-density regime Garzó, Brito, and Soto (2024a, b) shows that the effect of density on the phase diagrams for the marginal segregation curve ($\Lambda=0$) is generally significant. Specifically, the RBNE region ($\Lambda<0$) is increased as the system becomes more dense in the absence of gravity (see Fig. 11), but decreases in the presence of gravity (see Figs. 12 and 13).
Since the results derived in the paper have been obtained from the (inelastic) Enskog kinetic equation, one could speculate about the range of validity for reproducing the results obtained from MD simulations. As for molecular (elastic) mixtures, the Enskog equation provides a semiquantitative description of the hard-sphere system that neglects velocity correlations between the particles that are about to collide (molecular chaos hypothesis). Although the existence of these velocity correlations restricts the range of validity of the (inelastic) Enskog theory, the latter can be still considered as a good approximation (especially at the level of the transport properties) for both moderate inelasticities and/or densities. In particular, the Enskog results have been shown in general to compare quite well with MD simulationsLutsko, Brey, and Dufty (2002); Mitrano et al. (2011); Chialvo and Sundaresan (2013); Mitrano, Garzó, and Hrenya (2014) and even with real experimentsYang et al. (2002); Huan et al. (2004) for moderately high densities such as $n\sigma^{3}\lesssim 0.25$ for $d=3$. In the case of the $\Delta$-model, Fig. 2 highlights the good agreement for the shear viscosity over the entire range of the coefficient of normal restitution, even at relatively high density ($\phi=0.314$). Thus, within the context of the $\Delta$-model, we expect that the range of validity of the present results to be similar to that found in previous works for the conventional IHS model.
It is important to recall that our theory applies to relatively dilute granular materials, in which grains mostly interact through collisions rather than enduring contacts. This means that since grains move quickly, inertial effects become important. To measure the influence of inertia on the dynamic properties of grains, the inertial number $P$ is usually introduced. For a monocomponent granular gas, $P=\dot{\gamma}\sigma/\sqrt{p/\rho_{m}}$, where $\dot{\gamma}$ is the shear rate and $\rho_{m}$ is the particle density. Since our results are restricted to the Navier–Stokes domain (where $\dot{\gamma}$ is very small), the transport coefficients are independent of the shear rate or the inertia number. An interesting problem is extending our results to shearing granular flows to determine the nonlinear dependence of the rheological properties on the shear rate, or equivalently, on the inertia number. This will enable us to compare our results with those of previous studies on polyatomic moleculesGrmela and Lafleur (1998); Eu and Ohr (2001) and with those obtained using micropolar generalized Navier–Stokes equations.Eringen (1966)
@@
As said before, another future challenging problem within the $\Delta$-model is to determine the heat flux transport coefficients for dense mixtures. Knowing the complete set of Navier-Stokes transport coefficients will enable us to perform a linear stability analysis of the homogeneous steady state, among other applications. In particular, since the homogeneous steady state is stable in the dilute limit Garzó, Brito, and Soto (2021), we want to see if density corrections to the transport coefficients can modify the stability of the homogeneous state. Additionally, the reliability of the theoretical results derived here, which were obtained under certain approximations, should be assessed against computer simulations. One possible project is to carry out computer simulations to measure shear viscosity. As for dry granular mixtures, Montanero and Garzó (2003); Garzó and Montanero (2003); Chamorro and Garzó (2023) we plan to perform simulations on a granular mixture subjected to simple shear flow, where the cooling effects associated with collision dissipation are precisely offset by the presence of the $\Delta$ parameters in the collisional rules. Work along these lines is in progress.